# A Survey on 3D Gaussian Splatting

### v2

@@ 6.4 Performance Benchmarking: Rendering Quality in Driving Scenes
$\bullet$ Dataset: nuScenes [204] dataset is a comprehensive collection for autonomous driving research, featuring 1000 driving scenes captured using an array of sensors including six cameras, one LiDAR, five RADARs, GPS, and IMU. It provides detailed annotations for 23 object classes with 3D bounding boxes. Six challenging scenes are used for evaluation [213].
@@
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve one recent paper which models driving scene with 3D GS [213], as well as six NeRF based approaches [234, 10, 11, 235, 236, 237].
$\bullet$ Result: The results in Table IV demonstrate that the 3D GS based methods significantly surpass the NeRF based methods across all evaluated metrics. For instance, DrivingGaussian-L [213] outperforms S-NeRF [236] by 3.31dB in terms of PSNR. This suggests that 3D Gaussians can benefit from multi-sensor information to capture dynamic objects in driving scenes, especially fast-moving dynamic objects.
6.5 Performance Benchmarking: Human Avatar
@@
TABLE VI: Quantitative surgical 3D reconstruction results6.6) on EndoNeRF [216], in terms of PSNR, SSIM, and LPIPS. The numbers of non-GS methods and FPS are taken from [220]. denotes the average of the values reported in the original paper.
@@
Endo-4DGS [219] [arXiv] 37.00 0.96 0.05 - 4GB
@@
EndoGS [221] [arXiv] 37.94 0.97 0.03 - -
6.6 Performance Benchmarking: Surgical 3D Reconstruction
$\bullet$ Result: Table VI shows that introducing the explicit representation of 3D Gaussians leads to several significant improvements. For instance, EndoGaussian [220] outperforms a strong baseline, LerPlane-32k [217], among all metrics. In particular, EndoGaussian demonstrates an approximate 224-fold increase in speed while consumes just 10% of the GPU resources. These impressive results attest to the efficiency of GS-based methods, which not only expedite processing but also minimize GPU load, thus easing the demands on hardware. Such attributes are vitally significant for real-world surgical application deployment, where optimized resource usage can be a crucial determinant of practical utility.
@@ 7 Future Research Directions
$\bullet$ 3D GS for Robotics. In the world of robotics, especially when it comes to robots performing tasks that involve handling objects in ways similar to humans, there is a growing need for these machines to be able to navigate and manipulate their environment in a more intuitive and dynamic manner. This necessity stems from the desire to deploy intelligent robots in real-world settings, where they are often faced with new and unfamiliar tasks. Traditional approaches to robotic manipulation have relied heavily on understanding the environment through semantic representation, which means identifying objects and their attributes. However, these methods often overlook the importance of how things move and interact over time, which is crucial for completing tasks in the way humans intend. Imagine a robot trying to stack blocks based on a verbal command. Using only semantic representation, the robot might recognize the blocks but fail to stack them if it does not understand how the blocks should be positioned and moved relative to each other over time. This gap in understanding can lead to failure in executing tasks that require interaction with multiple objects or navigating dynamic environments. Due to its explicit nature, 3D GS can be used beyond mere semantic and structural analysis of environments; it also encompasses the dynamic aspect, providing a comprehensive understanding of how scenes evolve and objects interact over time. This is essential for robots tasked with navigating and manipulating their surroundings. Although initial efforts [246, 247] to construct world models using 3D GS are promising, they represent only the beginning of what is possible. Further research in this area is anticipated to enhance the capabilities of robots in performing tasks that require an understanding of both the physical space and the temporal changes within it.
@@
$\bullet$ 3D Large-scale Scene Reconstruction. 3D scene reconstruction and rendering, particularly for large-scale environments, has garnered considerable interest due to its applications in various fields, such as AR, VR, and autonomous driving. It is impractical to use 3D GS directly for large-scale scene reconstruction due to: i) reconstructing large-scale scenes within the current framework is unfeasible for the hardware; and ii) rendering so many Gaussians introduces a substantial computational load. As an early attempt, CityGS [248] proposes a divide-and-conquer training strategy alongside a level-of-detail rendering approach, by partitioning a large scene into smaller, manageable blocks, each represented by a reduced number of Gaussians. Overall, reconstructing large-scale scenes using 3D GS still holds significant potential waiting to be explored.
@@
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., camera pose estimation [249], the capture of hand-object interactions [250], the quantification of uncertainty [251], surgical scene reconstruction [219, 221, 220], and fluid synthesis [252]. These preliminary findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.
@@ 8 Conclusions
To the best of our knowledge, this survey presents the first comprehensive overview of 3D GS, a groundbreaking technique revolutionizing explicit radiance fields, computer graphics, and computer vision. It delineates the paradigm shift from traditional NeRF based methods, spotlighting the advantages of 3D GS in real-time rendering and enhanced editability. Our detailed analysis demonstrates the superiority of 3D GS in practical applications, particularly those highly sensitive to latency. We offer insights into principles, prospective research directions, and the unresolved challenges within this domain. Overall, 3D GS stands as a transformative technology, poised to significantly influence future advancements in 3D reconstruction and representation. This survey is intended to serve as a foundational resource, propelling further exploration and progress in this rapidly evolving field.

### v3

@@ 6.4 Performance Benchmarking: Driving Scenes
$\bullet$ Dataset: nuScenes [241] dataset is a comprehensive collection for autonomous driving, featuring 1000 driving scenes captured using an array of sensors including six cameras, one LiDAR, five RADARs, GPS, and IMU. It provides detailed annotations for 23 object classes with 3D bounding boxes. Six challenging scenes are used for evaluation [137].
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve one 3D GS based approach [137], as well as six NeRF based methods [266, 8, 9, 267, 268, 269].
$\bullet$ Result: The results in Table IV demonstrate that the 3D GS based methods significantly surpass the NeRF based methods across all evaluated metrics. For instance, DrivingGaussian-L [137] outperforms S-NeRF [268] by 3.31dB in terms of PSNR. This suggests that 3D Gaussians can benefit from multi-sensor information to capture dynamic objects in driving scenes, especially fast-moving ones.
6.5 Performance Benchmarking: Human Avatar
@@
6.6 Performance Benchmarking: Surgical Scenes
TABLE VI: Quantitative surgical 3D reconstruction results6.6) on EndoNeRF [250], in terms of PSNR, SSIM, and LPIPS. The numbers of non-GS methods, FPS, and GPU usage (Mem.) are taken from [141]. denotes numbers taken from [137]. denotes the average of the values reported in the original paper.
@@ 6.6 Performance Benchmarking: Surgical Scenes
EndoGS [142] [arXiv] 36.84 0.96 0.04 - -
Endo-4DGS [140] [MICCAI24] 37.00 0.96 0.05 - 4GB
@@
HFGS [143] [arXiv] 38.14 0.97 0.03 - -
$\bullet$ Result: Table VI shows that introducing the explicit representation of 3D Gaussians leads to several significant improvements. For instance, EndoGaussian [141] outperforms a strong baseline, LerPlane-32k [251], among all metrics. In particular, EndoGaussian demonstrates an approximate 224-fold increase in speed while consumes just 10% of the GPU resources. These impressive results attest to the efficiency of GS-based methods, which not only expedite processing but also minimize GPU load, thus easing the demands on hardware. Such attributes are vitally significant for real-world surgical application deployment, where optimized resource usage can be a key determinant of practical utility.
@@ 7 Future Research Directions
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these physical priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and realism of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the Spring-Mass 3D Gaussians for elastic object modeling [277], and the generalizable Gaussian representation based on multi-view stereo [278]. Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
$\bullet$ 3D GS for Robotics. In the world of robotics, especially when it comes to robots performing tasks that involve handling objects in ways similar to humans, there is a growing need for these machines to be able to navigate and manipulate their environment in a more intuitive and dynamic manner. This necessity stems from the desire to deploy intelligent robots in real-world settings, where they are often faced with new and unfamiliar tasks. Traditional approaches to robotic manipulation have relied heavily on understanding the environment through semantic representation, which means identifying objects and their attributes. However, these methods often overlook the importance of how things move and interact over time, which is crucial for completing tasks in the way humans intend. Imagine a robot trying to stack blocks based on a verbal command. Using only semantic representation, the robot might recognize the blocks but fail to stack them if it does not understand how the blocks should be positioned and moved relative to each other over time. This gap in understanding can lead to failure in executing tasks that require interaction with multiple objects or navigating dynamic environments. Due to its explicit nature, 3D GS can be used beyond mere semantic and structural analysis of environments; it also encompasses the dynamic aspect, providing a comprehensive understanding of how scenes evolve and objects interact over time. This is essential for robots tasked with navigating and manipulating their surroundings. Although initial GS based efforts on world models [279, 280] and reinforcement learning [281, 282] are promising, they represent only the beginning of what is possible. Further research in this area is anticipated to enhance the capabilities of robots in performing tasks that require an understanding of both the physical space and the temporal changes within it.
@@
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., the capture of hand-object interactions [285], the quantification of uncertainty [286], and fluid synthesis [287]. These preliminary findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.
@@ 8 Conclusions
To the best of our knowledge, this survey presents the first comprehensive overview of 3D GS, a groundbreaking technique revolutionizing explicit radiance fields, computer graphics, and computer vision. It delineates the paradigm shift from traditional NeRF based methods, spotlighting the advantages of 3D GS in real-time rendering and enhanced editability. Our in-depth analysis and extensive quantitative studies demonstrate the superiority of 3D GS in practical applications, particularly those highly sensitive to latency. We offer insights into principles, prospective research directions, and the unresolved challenges within this domain. Overall, 3D GS stands as a transformative technology, poised to significantly influence future advancements in 3D reconstruction and representation. This survey is intended to serve as a foundational resource, propelling further exploration and progress in this rapidly evolving field.